30 individuals were asked the following question: What is your favorite soft drink? The table below shows the results.

Names

How many times a week do you typically exercise?

1

Diet Coke

2

Diet Coke

3

Pepsi

4

Diet Pepsi

5

Sprite

6

Mountain Dew

7

Diet Coke

8

Coke Zero

9

Sprite

10

Fanta

11

Root Beer

12

Sierra Mist

13

Diet Coke

14

Diet Dr. Pepper

15

Pepsi

16

Diet Pepsi

17

Diet Coke

18

Root Beer

19

Diet Coke

20

Sierra Mist

21

Diet Coke

22

Mountain Dew

23

7 UP

24

Sprite

25

Diet Coke

26

Root Beer

27

Dr. Pepper

28

7UP

29

Diet Coke

30

Diet Coke

1. Calculate the sample proportion for individuals that answered “Diet Coke”. SHOW THE CALCULATION

2. Calculate the appropriate SE based on the 30 individuals and #1 above. SHOW THE CALCULATION.

3. Calculate a 95% confidence interval for the proportion associated with the individuals who answered “Diet Coke”. SHOW ALL CALCULATIONS.

4. Interpret the full meaning of the confidence interval you have in #3. Remember to include all four components to an interpretation.
Answer #1

1. The sample proportion for individuals that answered “Diet Coke” = 10/30 = 0.33

2. SE = √p(1-p)/n = √0.33(1-0.33)/30 = 0.086

3. The 95% confidence interval will be:

= p ± z*(√p(1-p)/n

= 0.33 ± 1.96*(√0.33(1-0.33)/30

= (0.1646, 0.502)

4. We are 95% confident that the true proportion associated with the individuals who answered “Diet Coke” is between 0.1646 and 0.502.

Please give me a thumbs-up if this helps you out. Thank you!